OpenStudy (anonymous):

The distance covered by an object is governed by the expression x3 + 3x2 + 2x. If the time taken to cover this distance is given by x + 1, what is the expression that governs the speed of the object? x2+2x x2+2x−1x+1 x+2 x2+1x+1

OpenStudy (anonymous):

what a weird way to ask you to find \[\frac{3^3+3x^2+2x}{x+1}\]

OpenStudy (anonymous):

\[\frac{x^3+3x^2+2x}{x+1}\]actually

OpenStudy (anonymous):

im at total loss I have no idea what this question is asking for ive been outta school 8 months

OpenStudy (anonymous):

they made a silly word problem is all distance divided by time is rate, so really they are asking you to divide one polynomial by another

OpenStudy (anonymous):

i wrote above exactly what it is asking in math rather than in words, namely it is asking for \[\frac{x^3+3x^2+2x}{x+1}\]

OpenStudy (anonymous):

I didn't learn anything about polynomials yet I don't know which one it would be

OpenStudy (anonymous):

my teach thinks the third one

OpenStudy (anonymous):

do you know how to divide polynomials?

OpenStudy (anonymous):

your teacher "thinks" it is the third one??!!

OpenStudy (anonymous):

(hint, it is definitely not the third one!)

OpenStudy (anonymous):

my teachers a history teacher and im stuck doing math all day n I don't know how to divide them at all

OpenStudy (anonymous):

this topic was new to me 5 mins ago

OpenStudy (anonymous):

it is almost impossible for me to show you how to divide polynomials here you can read about it if you like on the other hand, if you just want the answer, once you know that the question is really a division problem, and since you are on the computer, use the computer to solve it

OpenStudy (anonymous):

i will send a link with the answer you can see what i wrote, and where the answer is

OpenStudy (anonymous):

thank you very much bc I tried everything and im struggling

OpenStudy (anonymous):

look at "alternate forms" you will find \(x(x+2)\) which is not an answer choice, but it is the same as \[x^2+2x\]

OpenStudy (anonymous):

that website is there a fee to get calcs on it like that

OpenStudy (anonymous):

you can pay if you want special features, but no, it is free also very smart, so be careful how you use it

OpenStudy (anonymous):

notice the use of parentheses when i made the fractions

OpenStudy (anonymous):

noted I appreciate your help

OpenStudy (anonymous):

no problem